OpenStudy (anonymous):

whaat is the slope of the line that connects the ordered pairs (9,-4) and (17,-36)

OpenStudy (zpupster):

slope is calculated as \[m=\frac{ (y_2 - y_1) }{ x_2 - x_1}\] so your (x,y )pairs are \[(x_1,y_1) = (9,-4)\ and \ (x_2,y_2) =(17,-36)\] so plug in your values \[\frac{ (-36-(-4)) }{ 17-9 }\]

OpenStudy (anonymous):

thank you